# Silicon Photomultipliers for Detection of Photon Bunching Signatures

### v1

@@ Abstract
In this work, photon bunching from LED light was observed for the first time using SiPMs. The bunching signature was observed with a significance of $7.3~{}\sigma$ using 97 hs of data. The light was spectrally filtered using a 1 nm bandpass filter and an Etalon filter to ensure temporal coherence of the field and its coherence time was measured to be $\tau_{C}=(13.0\pm 1.3)$ ps. The impact of SiPM non-idealities in these kinds of measurements is explored, and we describe the methodology to process SiPM analog waveforms and the event selection used to mitigate these non-idealities.
@@ 1 Introduction
From biomolecules to astronomical bodies, from LIDAR to Deep Space Optical communications, photon detection is a tool with wide range of scientific and technological applications. Thus, the development of new sensors enables new discoveries and pushes the frontier of knowledge in several disciplines. The ultimate photo-detector should be capable of resolving single photons with a high dynamic range, high detection efficiency and short response time, have photon-number resolution and operate at room temperature. In recent years, a novel device called Silicon Photomultiplier (SiPM) [1, 2] was developed that, in principle, fulfills all these requirements. Currently, SiPMs have been successfully implemented in medical imaging [3, 4], particle physics detectors [5, 6, 7, 8], astrophysics [9, 10] and communications applications [11, 12], among others. Nevertheless, applications to Quantum Optics and Photonics have been hindered in the past due to some of the SiPM’s non-idealities, which can be classified in two: dark counts and correlated noise. Dark counts are generated due to thermal excitation of electrons in the Silicon and so they are generated even if no light is incident on the sensors. Correlated noise has two distinct sources: crosstalk and afterpulsing. Crosstalk events occur when a primary dark or photon count produces a prompt secondary avalanche in the Silicon. Such events have an amplitude of two detected photons. In turn, afterpulsing are events in which, after a primary dark or photon count, a delayed avalanche is generated, producing two distinct events. These effects are called correlated noise because they require a primary count to appear and they degrade the sensor performance. Recently, the impact of these features in the field of quantum optics has been studied [13].
@@
One of the landmark experiments that gave rise to the field of quantum optics was the measurement of intensity fluctuations of light detected from thermal sources, first done by Hanbury-Brown and Twiss [14]. They observed photon correlations by measuring the arrival time difference between them at two separate detectors. These correlations are referred to as HBT effect or photon bunching. Besides opening a new field of research and giving an insight into the nature of light, Hanbury-Brown and Twiss applied this technique to astronomy and used it to measure the angular diameters of several stars [15].
@@
Nowadays, photon correlation measurements are applied to several fields, like time-correlated photon detection in time-domain diffuse optics [16], quantum version of Time-of-Flight LIDAR [17] and quantum image scanning microscopy [18]. In addition, advances in instrumentation technology have given new life to the application of the HBT technique in astrophysics. For example, the VERITAS [19] and MAGIC [20] collaborations have measured the angular diameter of various stars, and prospects for Cherenkov Telescope Array (CTA) were studied [21]. Moreover, the use of the technique was recently proposed for measuring gravitational waves [22].
@@ 2 Signal-to-Noise Ratio
The usual sensors used in Quantum Optics are the avalanche photodiodes (APDs). The traditional way to measure photon bunching is to use a beamsplitter and two sensors that are triggered in coincidence [23, 24] and the time difference between photon arrivals, $dt$, is measured. In this approach, a Signal-to-Noise Ratio (SNR) can be defined as the ratio between bunched events to accidental coincidence fluctuations in a resolving time $\tau_{R}$ as [15]
@@
where $\tau_{c}$ is the coherence time of the field, $R_{i}$ is the photon rate in detector $i$, $V$ is the degree of spatial coherence, $T$ is the integration time and $\tau_{R}$ is the resolving time of the experiment. This SNR formula is valid for linearly polarized light. A time-difference histogram will have a background of accidental events and a bunching signature, centered at zero time-difference if the optical path of both sensors is the same. The background fluctuations are accidental (random) coincidences of sensor counts, which mainly come from non-correlated light. In a case where the sensor has non-negligible dark counts, like in an SiPM, these have to be taken into account. If the resolving time $\tau_{R}$ is much larger than the coherence time $\tau_{c}$, the rate of coincidences $R_{c}$ in a window of duration $\tau_{R}$ is [25, 26]
$R_{c}=\overbrace{(R_{1}R_{2}+D_{1}D_{2})\tau_{R}}^{\text{accidental}}+\overbrace{R_{1}R_{2}\tau_{c}V^{2}}^{\text{signal}}\ ,$ (2)
where $D_{i}$ is the Dark Count Rate (DCR) in detector $i$. From equation (2), it is possible to obtain a relation between the SNR using APDs with respect to using SiPMs as
$\mathrm{SNR}_{\mathrm{SiPM}}=\frac{1}{\sqrt{F}}\mathrm{SNR}_{\mathrm{APD}}\ .$ (3)
Above we defined the factor F as $F=1+\frac{D_{1}D_{2}}{R_{1}R_{2}}$, and we assumed negligible dark counts for the APD case. Equation (3) shows the impact of this non-ideality of SiPMs sensors in this kind of measurement. In addition, the ratio of bunching events to accidental events in a time $\tau_{R}$, which is calculated as $C=\frac{\tau_{c}}{\tau_{R}}V^{2}$ [25], will also decrease when comparing the APD and the SiPM cases:
$C_{\mathrm{SiPM}}=\frac{1}{F}\frac{\tau_{c}}{\tau_{R}}V^{2}=\frac{1}{F}C_{\mathrm{APD}}\ \ .$ (4)
It can be seen that $C_{\mathrm{SiPM}}$ and $\mathrm{SNR}_{\mathrm{SiPM}}$ approach the APD case when $F\sim 1$.
@@ 3 Experimental Setup
The experimental setup used to measure photon bunching using SiPMs is shown in Figure 1. The setup were mounted inside a light tight facility specifically designed for SiPM characterization studies [27, 28].
[…]
The light source used for the experiment was a fiber coupled LED from THOR LABS (TL M530F2) with a center wavelength of 530 nm. Its output was fed into a single mode optical fiber, which is used to enforce spatial coherence of the field. A rate of $10^{10}$ photons per second was obtained at the output of the fiber collimator. A linear polarizer was placed in the beam path. Then, to increase the coherence time of the field, light was spectrally filtered using a 1 nm FWHM filter (TL FL05532-1) and an etalon filter (Light Machinery, OP-7423-1686-1). The etalon free spectral range is such that approximately 17 peaks fall in the 1 nm FWHM optical passband of the interference filter [29]. The coherence time of this configuration was estimated to be of the order of 10 ps [25]. Finally, the beam was separated with 50:50 beamsplitter and both SiPMs were placed at the same distance from it. Each SiPM (ONSEMI MicroFC-10035-SMT) was connected to an Analog Front-End board designed in-house with a time resolution below 1 ns. SiPMs were biased at 5 V overvoltage using a SourceMeter Unit and the SiPM DC current was monitored during the whole experiment. No temperature control on them was performed, but the experiment room was kept at $(22\pm 1)$ ${}^{\mathrm{o}}$. SiPMs were both placed at a distance of $(85\pm 1)$ cm from the collimator output. SiPM analog waveforms were acquired using a digitizer at 500 MSps (PicoScope 2406B) and the coincidence window used was 100 ns. A 3 ns cable delay was placed in one of the SiPM channels to help rule out hardware glitches.
Two experiments were performed: First, a signal measurement where the field was both spatially and temporally coherent (and thus bunching could be observed). Then, a background measurement where the effect was not present, because the field was purposely not temporally coherent. To remove the temporal coherence of the field, the 1 nm filter and the Etalon filter were removed and replaced with an optical attenuator to avoid SiPM signal saturation. In the case where photon bunching is not present, the background is expected to be uniform at short times.
@@
The measured coincidence rate between the two detectors in the 100 ns coincidence gate was about 10 kcps. Considering the acquisition dead time, both signal and background experiments were run for an effective time of 97 hs. Using equation (3), the expected SNR for that effective acquisition time is 96.
@@ 4 Event Selection
    No trigger timestamps present inside the baseline calculation samples.
@@ 5 Results and Discussion
After event selection, the time-difference histogram was constructed for the two experiments performed. On both histograms, an unwanted sinusoidal-like systematic pattern with a period of 2 ns was observed. The reason for this effect is that the digitizer-board channels are only synchronized down to its sampling period of 2 ns. The top histogram of Figure 3 shows the raw data points with this structure. The amplitude of this electronic jitter is negligible compared to the mean entry number, but relevant for observing the bunching peak. Due to this, a procedure had to be devised to remove this systematic effect. First, the time-difference histogram of the background dataset was cut into 2 ns slices. Then, all these slices were folded together and averaged into a single 2-ns window, where random noise was mitigated. This 2 ns window was then repeated along the time-difference axis to build a template of the systematic structure. This template was then used to normalize the raw histogram, and it is shown on the top plot of Figure 3. On the bottom plot of that same figure, the normalized histogram is shown. A uniform fit was performed on the normalized background with a resulting $\chi^{2}/\mathrm{dof}=99.7/100$ and a p-value of 0.49. In addition, a Kolmogorov-Smirnoff (KS) test [30] was run to compare the background to a uniform distribution, and a p-value of 0.37 was obtained. Both tests are consistent with the expected background distribution in the absence of bunching signatures.
Figure 3: Top plot: raw data histogram of timestamp differences in the background experiment, along with the generated template used for the removal of the systematic 2 ns structure. Bottom plot: normalized timestamp difference histogram in the background experiment and a uniform fit with resulting $\chi^{2}/\mathrm{dof}=99.7/100$ and a p-value of 0.49.
@@
Figure 4: Top plot: raw data histogram of timestamp differences in the signal experiment, along with the generated template used for the removal of the systematic 2 ns structure. Bottom plot: normalized timestamp difference histogram of both experiments. The electrical delay of 3 ns can be seen in green.
@@
To further identify the characteristics of the peak, a Gaussian fit was performed on it. The coherence time of the field is much smaller that the overall time resolution of the detection system, which is a combination of SiPM and electronic jitter and the digital pulse-processing algorithms. This is why the shape of the bunching peak is expected to have a Gaussian distribution [31]. Figure 5 is a zoom of the time-difference histogram centered in the bunching peak, where a Gaussian fit was performed ($\chi^{2}/\mathrm{dof}=18/16$ and $\mathrm{p-value}=0.31$).
Figure 5: Signal peak with a Gaussian fit. The peak is expected to be Gaussian because it is dominated by the overall time resolution of the detection system. The fit has a $\chi^{2}/\mathrm{dof}=18/16$, which results in a $\mathrm{p-value}=0.31$. The width of the peak is $(920\pm 83)$ ps FWHM, the peak’s center is $(3.15\pm 0.03)$ ns and the peak height is $(3.3\pm 0.2)\cdot 10^{-3}$. The system resolving time, given by the composite jitter of the two SiPM detectors is shown in red (centered around the peak maximum).
The width of the Gaussian distribution obtained from the fit is $(920\pm 83)$ ps FWHM, which represents the overall time resolution of the system, or the resolving time $\tau_{R}$. The mean obtained from the fit is $(3.15\pm 0.03)$ ns, compatible with the $\sim 3$ ns cable delay introduced in one of the channels. The height obtained from the fit is $(3.3\pm 0.2)\cdot 10^{-3}$. From the latter, and using Eq. 4, the coherence time $\tau_{C}$ is $(13.0\pm 1.3)$ ps, which is consistent with the expected value of the order of 10 ps. The experimental SNR was calculated to be 99, using events in one resolving time. This value is consistent with the estimation given by equation 3.

### v3

# Biblio.bib

@@ Abstract
In this work, photon bunching from LED light was observed for the first time using SiPMs. The bunching signature was observed with a significance of $7.3~{}\sigma$ using 97 hs of data. The light was spectrally filtered using a 1 nm bandpass filter and an Etalon filter to ensure temporal coherence of the field and its coherence time was measured to be $\tau_{C}=(19\pm 2)$ ps. The impact of SiPM non-idealities in these kinds of measurements is explored, and we describe the methodology to process SiPM analog waveforms and the event selection used to mitigate these non-idealities.
@@ 1 Introduction
From biomolecules to astronomical bodies, from LIDAR to Deep Space Optical communications, photon detection is a tool with wide range of scientific and technological applications. Thus, the development of new sensors enables new discoveries and pushes the frontier of knowledge in several disciplines. The ultimate photo-detector should be capable of resolving single photons with a high dynamic range, high detection efficiency and short response time, have photon-number resolution and operate at room temperature. In recent years, a novel device called Silicon Photomultiplier (SiPM) [nepomuk_3_sipms, sipm_review] was developed that, in principle, fulfills all these requirements. Currently, SiPMs have been successfully implemented in medical imaging [pet1, pet2], particle physics detectors [hep1, hep2, hep3, hep4], astrophysics [ap1, ap2] and communications applications [comm1, comm2], among others. Nevertheless, applications to Quantum Optics and Photonics have been hindered in the past due to some of the SiPM’s non-idealities, which can be classified in two: dark counts and correlated noise. Dark counts are generated due to thermal excitation of electrons in the Silicon and so they are generated even if no light is incident on the sensors. Correlated noise has two distinct sources: crosstalk and afterpulsing. Crosstalk events occur when a primary dark or photon count produces a prompt secondary avalanche in the Silicon. Such events have an amplitude of two detected photons. In turn, afterpulsing are events in which, after a primary dark or photon count, a delayed avalanche is generated, producing two distinct events. These effects are called correlated noise because they require a primary count to appear and they degrade the sensor performance. Recently, the impact of these features in the field of quantum optics has been studied [chesi2019].
@@
One of the landmark experiments that gave rise to the field of quantum optics was the measurement of intensity fluctuations of light detected from thermal sources, first done by Hanbury-Brown and Twiss [hbt_1956]. They observed photon correlations by measuring the arrival time difference between them at two separate detectors. These correlations are referred to as HBT effect or photon bunching. The name “photon bunching” alludes to the fact that a fraction of the detected photon events have an increased probability of arriving together. Besides opening a new field of research and giving an insight into the nature of light, Hanbury-Brown and Twiss applied this technique to astronomy and used it to measure the angular diameters of several stars [hbt].
@@
Nowadays, photon correlation measurements are applied to several fields, like time-correlated photon detection in time-domain diffuse optics [dalla_mora], quantum version of Time-of-Flight LIDAR [tan_2023] and quantum image scanning microscopy [lubin_2019]. In addition, advances in instrumentation technology have given new life to the application of the HBT technique in astrophysics. For example, the VERITAS [abeysekara_2020] and MAGIC [delgado2021] collaborations have measured the angular diameter of various stars, and prospects for Cherenkov Telescope Array (CTA) were studied [dravins2012]. Moreover, the use of the technique was recently proposed for measuring gravitational waves [park2021].
@@ 2 Signal-to-Noise Ratio
The usual sensors used in Quantum Optics are the avalanche photodiodes (APDs). The traditional way to measure photon bunching is to use a beamsplitter and two sensors that are triggered in coincidence [tan_paper, tan_thesis] and then measure the time difference between photon arrivals, $dt$. In this approach, a Signal-to-Noise Ratio (SNR) can be defined as the ratio between bunched events to total coincidence fluctuations in a resolving time $\tau_{R}$. The resolving time is the overall time resolution of the experimental setup. The main contribution to this time is the 2-SiPM’s composite jitter. The total coincidence fluctuations are just the Poisson fluctuations of total counts (and so they can be calculated as $\sqrt{\mathrm{total\ counts}}$). These have two contributions: The accidental (non-bunched photons plus dark counts) event fluctuations and the signal (bunched photons) event fluctuations. It is usually the case that the accidental count fluctuations are much larger than the signal fluctuations, so these are often not considered in this calculation. With this in mind, the SNR can be calculated as [hbt]
@@
where $\tau_{c}$ is the coherence time of the field, $R_{i}$ is the photon rate in detector $i$, $V$ is the degree of spatial coherence, $T$ is the integration time and $\tau_{R}$ is the resolving time of the experiment. The coherence time is the time in which the phase of the field remains correlated or predictable, and the integration time is the total measurement time of the experiment. This SNR formula is valid for linearly polarized light, which will be the case for our experiment.
A time-difference histogram with the registered events will have a background of accidental events and a bunching excess peak, centered at zero time-difference if the optical path of both sensors is the same. When normalized using the accidental background [mandel], this histogram is the second order correlation function $g^{(2)}$. As previously mentioned, the background fluctuations are accidental (random) coincidences of sensor counts, which come from non-correlated light and dark counts. In a case where the sensor has non-negligible dark counts, like in an SiPM, these have to be taken into account. If the resolving time $\tau_{R}$ is much larger than the coherence time $\tau_{c}$, the rate of coincidences $R_{c}$ in a window of duration $\tau_{R}$ is [scarl_contrast]
$R_{c}$ $=\overbrace{N_{1}N_{2}\tau_{R}}^{\text{accidental}}+\overbrace{R_{1}R_{2}\tau_{c}V^{2}}^{\text{signal}}$ (2)
$=N_{1}N_{2}\tau_{R}\Big{(}1+\frac{R_{1}R_{2}}{N_{1}N_{2}}\frac{\tau_{c}}{\tau_{R}}V^{2}\Big{)}\ ,$
where $D_{i}$ is the Dark Count Rate (DCR) in detector $i$ and $N_{i}=R_{i}+D_{i}$ is the total rate in detector $i$. Note that the presence of bunching results in an excess of events over the accidental background $N_{1}N_{2}\tau_{R}$. From Equation (2), it is possible to obtain a relation between the SNR using APDs with respect to using SiPMs as
$\mathrm{SNR}_{\mathrm{SiPM}}=\frac{R_{1}R_{2}\tau_{c}V^{2}T}{\sqrt{N_{1}N_{2}\tau_{R}T}}=\frac{1}{\sqrt{F}}\mathrm{SNR}_{\mathrm{APD}}\ .$ (3)
Above we defined the factor $F=\frac{N_{1}N_{2}}{R_{1}R_{2}}$, and we assumed negligible dark counts for the APD case. To derive the above expression, the signal counts were divided by the fluctuation of accidental counts. Equation (3) shows the impact of this non-ideality of SiPMs sensors in these kinds of measurements. In addition, the ratio of bunching events to accidental events in a time $\tau_{R}$ [mandel] will also decrease when comparing the APD and the SiPM cases:
$C_{\mathrm{SiPM}}=\frac{R_{1}R_{2}}{N_{1}N_{2}}\frac{\tau_{c}}{\tau_{R}}V^{2}=\frac{1}{F}C_{\mathrm{APD}}\ \ .$ (4)
It can be seen that $C_{\mathrm{SiPM}}$ and $\mathrm{SNR}_{\mathrm{SiPM}}$ approach the APD case when $F\sim 1$ (or $R_{i}\gg D_{i}$).
@@ 3 Experimental Setup
The experimental setup used to measure photon bunching using SiPMs is shown in Figure 1. The setup was mounted inside a light tight facility specifically designed for SiPM characterization studies [spock1, spock2].
Figure 1: Schematic of the experimental setup used to detect photon bunching with SiPMs. A green LED source was coupled to a single-mode fiber to ensure spatial coherence of the field. Light was filtered with a 1 nm filter and an Etalon filter to ensure temporal coherence as well. Two 1 mm2 SiPMs were placed in both arms of a 50:50 beamsplitter to perform coincidence measurements.
The light source used for the experiment was a fiber coupled LED from THOR LABS (TL M530F2) with a center wavelength of 530 nm. Its output was fed into a single mode optical fiber, which is used to enforce spatial coherence of the field. A rate of $10^{10}$ photons per second was obtained at the output of the fiber collimator. A linear polarizer was placed in the beam path. Then, to increase the coherence time of the field, light was spectrally filtered using a 1 nm FWHM filter (TL FL05532-1) and an etalon filter (Light Machinery, OP-7423-1686-1). The etalon free spectral range is such that approximately 17 peaks fall in the 1 nm FWHM optical passband of the interference filter [etalon_spectrum]. The coherence time of this configuration was estimated to be of the order of 10 ps [mandel]. Finally, the beam was separated with 50:50 beamsplitter and both SiPMs were placed at the same distance from it. Each SiPM (ONSEMI MicroFC-10035-SMT) was connected to an Analog Front-End board designed in-house with a time resolution below 1 ns. SiPMs were biased at 5 V overvoltage using a SourceMeter Unit and the SiPM DC current was monitored during the whole experiment. No temperature control on them was performed, but the experiment room was kept at $(22\pm 1)$ oC. SiPMs were both placed at a distance of $(85\pm 1)$ cm from the collimator output. SiPM analog waveforms were acquired using a digitizer at 500 MSps (PicoScope 2406B). The DAQ is configured with a coincidence window of 100 ns, a record length of 200 ns and a pre trigger region of 95 ns. This means that at least one of the pulses in a captured event is approximately centered in the acquisition window. A 3 ns cable delay was placed in one of the SiPM channels to help rule out hardware glitches.
Two experiments were performed: First, a signal measurement where the field was both spatially and temporally coherent (and thus bunching could be observed). Then, a background measurement where the effect was not present, because the field was purposely not temporally coherent. To remove the temporal coherence of the field, the 1 nm filter and the Etalon filter were removed and replaced with an optical attenuator to avoid SiPM signal saturation. In the case where photon bunching is not present, the background is expected to be uniform at small time differences.
@@
For reference, these values would result in $F\sim 1.6$. Compared to the use of APDs, this results in an SNR and a bunching-peak relative height 20 % and 38 % lower, respectively.
The measured coincidence rate between the two detectors in the 100 ns coincidence gate was about 10 kcps. Considering the acquisition dead time, both signal and background experiments were run for an effective time of 97 hs. Using Equation (3), the expected SNR for that effective acquisition time is 80, while it would be 100 in the APD case.
@@ 4 Event Selection
    No samples above the threshold on the baseline calculating region. This selection cut is employed because the baseline can’t be correctly estimated if a previous pulse is present in that region.
@@ 5 Results and Discussion
After event selection, the time-difference histogram was constructed for the two experiments performed. On both histograms, an unwanted sinusoidal-like systematic pattern with a period of 2 ns was observed. The reason for this effect is that the digitizer-board channels are only synchronized down to its sampling period of 2 ns. The top histogram of Figure 3 shows the raw data points with this structure. The amplitude of this electronic jitter is negligible compared to the mean entry number, but relevant for observing the bunching peak. Due to this, a procedure had to be devised to remove this systematic effect. First, the time-difference histogram of the background dataset was cut into 2-ns slices. Then, all these slices were averaged into a single 2-ns window. Only systematic artifacts survive this procedure, so this is a good method to estimate this structure generated from the imperfect DAQ hardware. This averaged 2-ns window was then repeated along the time-difference axis to build a template of the systematic structure. This template was then used to normalize the raw histogram, and it is shown on the top plot of Figure 3. On the bottom plot of that same figure, the normalized histogram is shown. A uniform fit was performed on the normalized background with a resulting $\chi^{2}/\mathrm{dof}=99.7/100$ and a p-value of 0.49. In addition, a Kolmogorov-Smirnoff (KS) test [ks_test] was run to compare the background to a uniform distribution, and a p-value of 0.37 was obtained. Both tests are consistent with the expected background distribution in the absence of bunching signatures.
Figure 3: Top plot: raw data histogram of timestamp differences in the background experiment, along with the generated template used for the removal of the systematic 2-ns structure. Bottom plot: normalized timestamp difference histogram in the background experiment and a uniform fit with resulting $\chi^{2}/\mathrm{dof}=99.7/100$ and a p-value of 0.49.
@@
Figure 4: Top plot: raw data histogram of timestamp differences in the signal experiment, along with the generated template used for the removal of the systematic 2-ns structure. Bottom plot: normalized timestamp difference histogram of both experiments. The electrical delay of 3 ns can be seen in green.
@@
To further identify the characteristics of the peak, a Gaussian fit was performed on it. The coherence time of the field is much smaller that the overall time resolution of the detection system, which is a combination of SiPM and electronic jitter and the digital pulse-processing algorithms. This is the reason why the shape of the bunching peak is expected to have a Gaussian distribution [sipm_jitter_1]. Figure 5 is a zoom of the time-difference histogram centered in the bunching peak, where a Gaussian fit was performed ($\chi^{2}/\mathrm{dof}=18/16$ and p-value of 0.31).
Figure 5: Signal peak with a Gaussian fit. The peak is expected to be Gaussian because it is dominated by the overall time resolution of the detection system. The fit has a $\chi^{2}/\mathrm{dof}=18/16$, which results in a p-value of 0.31. The width of the peak is $(920\pm 83)$ ps FWHM, the peak’s center is $(3.15\pm 0.03)$ ns and the peak height is $(3.3\pm 0.2)\cdot 10^{-3}$. The system resolving time, given by the composite jitter of the two SiPM detectors is shown in red (centered around the peak maximum).
The width of the Gaussian distribution obtained from the fit is $(920\pm 83)$ ps FWHM, which represents the overall time resolution of the system, or the resolving time $\tau_{R}$. The mean obtained from the fit is $(3.15\pm 0.03)$ ns, compatible with the $\sim 3$ ns cable delay introduced in one of the channels. The height obtained from the fit is $(3.3\pm 0.2)\cdot 10^{-3}$. From the latter, and using Equation (4), the coherence time $\tau_{c}=(19\pm 2)$ ps was obtained. This value is consistent with the expected value in the order of 10 ps. The experimental SNR was calculated to be 99 using events in one resolving time. This value is of the same order of magnitude as the preliminary estimation given with Equation (3).